# The global communication pathways of the human brain transcend the cortical - subcortical - cerebellar division

### v2

@@ Results
Figure 3: Contribution of anatomical components to cross-modular connectivity. a, b, Visualization of the structural connectivity at the level of anatomical components comprising the cortex, the cerebellum and nine subcortical structures; a, for the brain-wide network and b, the subnetworks representing network communities C1 to C4. Community C5 is omitted for its simplicity as it only contains three regions. Node diameters are proportional to the number of ROIs to each anatomical component. Link widths reflect the number of ROIs connected within the two components. c, Illustration of different roles that nodes can play within a modular and hierarchical network, and their characterization. Nodes can be either locally connected inside a module (yellow dot), be well rooted in a community but share connections to other communities (purple triangle), or uniformly participate over all communities (dark purple diamond). These differential contributions are typically characterized by two parameters allowing to project each node into a 2D map (axes on the right panel). (i) The degree, informing of the overall centrality of a node and (ii) the participation index, indicating how a node spreads its links among communities. Participation 0.0 means the node is only locally connected in one module (e.g., the blue dot) while participation 1.0 means a node is equally likely connected to all communities (e.g., the purple diamond). Locally connected nodes are expected to play more specialized functions while hubs with large participation are usually associated with multifunctional roles due to their access to a wider range of information. d, Mapping of the roles of all the ROIs, considering the network is divided into cortex, subcortex and cerebellum. For visual clarity, the 718 ROIs are presented in three separate maps for cortical, subcortical and cerebellar regions. The solid horizontal lines represent the average node degree and gray shadows the range for one standard deviation. e, Mapping of the roles for all the ROIs, considering the network is divided into the five network communities. Top left panel shows all the 718 ROIs on a single map. The remaining panels show the same results but individually for the ROIs of each of the 11 anatomical components studied. ROIs are colored according to the network community they were assigned to.
@@ Methods
The cortex and cerebellum are contra-laterally connected via the cortico-ponto-cerebellar and the cerebello-thalamo-cortical pathways Ramnani (2006); Palesi et al. (2015, 2017); D’Angelo (2018). To only display direct anatomical connections, ($i$) cortico-cerebellar tracts were filtered as diffusion imaging cannot distinguish between direct and indirect pathways. This resulted in the filtering of 8.3% of initial connections. ($ii$) Ipsilateral tracts between the cerebellum and the brainstem (pons), ($iii$) ipsilateral tracts between cerebellum and thalamus (total of 2.0% of initial connections) where filtered out, as well as ($iv$) direct inter-hemispheric cerebellar tracts (2.6% of initial tracts), as the cerebellar hemispheres are indirectly connected via the vermis D’Angelo (2018).
@@
Here, we define integration as the sensitivity of one node (or of a set of selected nodes) to external stimuli applied to the network Zamora-López et al. (2010). Let $X$ be the set of all nodes in a network and $H\subset X$ a subset of arbitrarily chosen nodes. Then, the integration capacity $I(H)$ is defined as the joint response by the nodes in subset $H$, conditional to the simultaneous stimuli applied at all other nodes, the complementary set $X-H$. Formally,
@@
For the results in Fig. 5b, $H$ is considered as the set of 30 ROIs with largest degree $k$, and $X-H$ are the remaining 688 ROIs. The hub sets $H$ are always ranked based on their global degree, as in Fig. 4a, but the selection for the different cases are based on the community the ROIs belong to, in either the anatomical or the network-based division. Hence, the colored bars all correspond to the integration capacity of 30 hubs, but only for those that belong to a specific module. The results in Fig. 5c are computed equivalently but considering hub sets $H$ of increasing size, from $|H|=1$ to $|H|=170$, adding one hub at a time to the set $H$.
@@ Acknowledgements.
We thank Egidio d’Angelo, Claudia Gandini Wheeler-Kingshott and Fulvia Palesi for the insightful and inspiring recommendations and discussions. This research was supported (G.Z.L.) by the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreement No. 945539 (Human Brain Project SGA3), (X.K.) by a short-term fellowship of the European Molecular Biology Organization (Grant No. 7366). J.S. is a FI fellow with the support of AGAUR, Generalitat de Catalunya and Fondo Social Europeo (2025 FI-2 00170). G.D. received funding from Grant PID2022-136216NB-I00 funded by MICIU/AEI/10.13039/501100011033, by ”ERDF A way of making Europe”, ERDF, EU, and by Project NEurological MEchanismS of Injury, and Sleep-like cellular dynamics (NEMESIS) (ref. 101071900) funded by the EU ERC Synergy Horizon Europe.

### v3

@@ Results
Figure 3: Contribution of anatomical components to cross-modular connectivity. a, b, Visualization of the structural connectivity at the level of anatomical components comprising the cortex, the cerebellum and nine subcortical structures; a, for the brain-wide network and b, the subnetworks representing network communities C1 to C4. Community C5 is omitted for its simplicity as it only contains three regions. Node diameters are proportional to the number of ROIs to each anatomical component. Link widths reflect the number of ROIs connected within the two components. c, Illustration of different roles that nodes can play within a modular and hierarchical network, and their characterization. Nodes can be either locally connected inside a module (yellow dot), be well rooted in a community but share connections to other communities (purple triangle), or uniformly participate over all communities (dark purple diamond). These differential contributions are typically characterized by two parameters allowing to project each node into a 2D map (axes on the right panel). (i) The degree, informing of the overall centrality of a node and (ii) the participation index, indicating how a node spreads its links among communities. Participation 0.0 means the node is only locally connected in one module (e.g., the blue dot) while participation 1.0 means a node is equally likely connected to all communities (e.g., the purple diamond). Locally connected nodes are expected to play more specialized functions while hubs with large participation are usually associated with multifunctional roles due to their access to a wider range of information. d, Mapping of the roles of all the ROIs, considering the network is divided into cortex, subcortex and cerebellum. For visual clarity, the 718 ROIs are presented in three separate maps for cortical, subcortical and cerebellar regions. The solid horizontal lines represent the average node degree and gray shadows the range for one standard deviation. e, Mapping of the roles for all the ROIs, considering the network is divided into the five network communities. Top left panel shows all the 718 ROIs on a single map. The remaining panels show the same results but individually for the ROIs of each of the $1$ anatomical components studied. ROIs are colored according to the network community they were assigned to.
@@ Methods
The cortex and cerebellum are contra-laterally connected via the cortico-ponto-cerebellar and the cerebello-thalamo-cortical pathways Ramnani (2006); Palesi et al. (2015, 2017); D’Angelo (2018). To only display direct anatomical connections, ($i$) cortico-cerebellar tracts were filtered as diffusion imaging cannot distinguish between direct and indirect pathways. This resulted in the filtering of 8.3% of initial connections. ($ii$) Ipsilateral tracts between the cerebellum and the brainstem (pons), (iiiiii) ipsilateral tracts between cerebellum and thalamus (total of 2.0% of initial connections) where filtered out, as well as ($iv$) direct inter-hemispheric cerebellar tracts (2.6% of initial tracts), as the cerebellar hemispheres are indirectly connected via the vermis D’Angelo (2018).
@@
Here, we define integration as the sensitivity of one node (or of a set of selected nodes) to external stimuli applied to the network Zamora-López et al. (2010). Let $X$ be the set of all nodes in a network and $H\subset X$ a subset of arbitrarily chosen nodes. Then, the integration capacity $I(H)$ is defined as the joint response by the nodes in subset $H$, conditional to the simultaneous stimuli applied at all other nodes, the complementary set XHX-H. Formally,
@@
For the results in Fig. 5b, $H$ is considered as the set of 30 ROIs with largest degree $k$, and XHX-H are the remaining 688 ROIs. The hub sets $H$ are always ranked based on their global degree, as in Fig. 4a, but the selection for the different cases are based on the community the ROIs belong to, in either the anatomical or the network-based division. Hence, the colored bars all correspond to the integration capacity of 30 hubs, but only for those that belong to a specific module. The results in Fig. 5c are computed equivalently but considering hub sets $H$ of increasing size, from $|H|=1$ to $|H|=170$, adding one hub at a time to the set $H$.
@@ Acknowledgements.
We thank Egidio d’Angelo, Claudia Gandini Wheeler-Kingshott and Fulvia Palesi for the insightful and inspiring recommendations and discussions. J.S. is a FI fellow of AGAUR, Generalitat de Catalunya and Fondo Social Europeo (2025 FI-2 00170). X.K. received support by a short-term fellowship of the European Molecular Biology Organization (Grant No. 7366). G.D. received funding from Grant PID2022-136216NB-I00 funded by MICIU/AEI/10.13039/501100011033, by ”ERDF A way of making Europe”, ERDF, EU, and by Project NEurological MEchanismS of Injury, and Sleep-like cellular dynamics (NEMESIS) (ref. 101071900) funded by the EU ERC Synergy Horizon Europe. G.Z.L. was supported by the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreement No. 945539 (Human Brain Project SGA3), and by ERDF-Project Brain dynamics, No. CZ.02.01.01/00/22-008/0004643.